# A Study of Optimizations for Fine-tuning Large Language Models

### v1

@@ 6 Conclusion and Future Work
In this paper, we analyzed state-of-the-art fine-tuning optimizations and their impact on memory and runtime during fine-tuning. Our results show that ZeRO-2 + LoRA is a reliable optimization default that typically provides the best balance between memory usage and fine-tuning runtime. Using the right set of optimizations, such as ZeRO-3 + LoRA + GC, we successfully fine-tuned models as large as Falcon-180B. Recognizing the need for fine-tuning LLMs under GPU constrained settings, we introduced an optimization matrix across model-size $\times$ context length. The optimization matrix can guide the user to select the right optimization mixture that enables fine-tuning for a given use case. We conclude that optimizations such as DeepSpeed ZeRO are essential for fine-tuning multi-billion parameter models as they support data and model parallelism, along with CPU off-loading. In conjunction with DeepSpeed ZeRO, it becomes necessary to select additional optimizations to avoid out-of-memory failures while balancing fine-tuning runtime.

### v2

@@ 6 Conclusion and Future Work
In this paper, we analyzed state-of-the-art fine-tuning optimizations and their impact on memory and runtime during fine-tuning. Our results show that ZeRO-2 + LoRA is a reliable optimization default that typically provides the best balance between memory usage and fine-tuning runtime. Using the right set of optimizations, such as ZeRO-3 + LoRA + GC, we successfully fine-tuned models as large as Falcon 180B. Recognizing the need for fine-tuning LLMs under GPU constrained settings, we introduced an optimization matrix across model-size $\times$ context length. The optimization matrix can guide the user to select the right optimization mixture that enables fine-tuning for a given use case. We conclude that optimizations such as DeepSpeed’s ZeRO are essential for fine-tuning multi-billion parameter models as they support data and model parallelism, along with CPU off-loading. In conjunction with ZeRO, it becomes necessary to select additional optimizations to avoid out-of-memory failures while balancing fine-tuning runtime.